If there are x teams in a league and each team plays every other team twice, the the total number of games (N) to be played is given by N=x^2-x. A chess league has 9 teams. What is the total number of games to be played if each team plays every other team twice?
Answer by stanbon(75887)

If there are x teams in a league and each team plays every other team twice, the the total number of games (N) to be played is given by N=x^2-x. A chess league has 9 teams. What is the total number of games to be played if each team plays every other team twice?
----------
N(x) = x^2-x
-----
Your Problem:
N(9) = 9^2-9 = 72
=======================
Cheers,
Stan H.